# Hierarchical Light Transformer Ensembles for Multimodal Trajectory Forecasting

### v1

@@ 4.2 Hierarchical Multimode Loss
Many existing MDN architectures rely on explicit clustering techniques, which can introduce challenges regarding robustness to changes in mode number. In contrast, our proposed approach embraces the hierarchical structure inherent in MDNs, allowing for implicit clustering. To integrate this hierarchical structure into our prediction mixture, we introduce a novel loss function aimed at softly enforcing proximity among modes belonging to the same meta-mode, termed the Hierarchical Winner-Takes-All (HWTA) loss. The HWTA loss, denoted as $\mathcal{L}$, comprises two distinct components: the meta-mode loss ($\mathcal{L}_{\text{meta}}$) and the Meta-Winner-Takes-All (MWTA) loss.
@@
with a hyperparameter $\gamma\in\left[0,1\right]$ to control the importance of both terms. We conduct a sensitivity study on this parameter in Appendix 0.D. The meta-mode loss is the NLL of the meta-mixture, considering $N\in\mathbb{N}$ as the number of meta-modes:
$\mathcal{L}_{\text{meta}}=\frac{1}{|D|}\sum^{|D|}_{i=1}\sum^{K^{\star}}_{k^{\star}=1}-\text{log}\left(\pi_{\theta_{k^{\star}}}(X_{i})\times\mbox{Laplace}(\overline{\boldsymbol{\mu}}_{\theta_{k^{\star}}}(X_{i}),\overline{\mathbf{b}}_{\theta_{k^{\star}}}(X_{i}))\right)$ (9)
This term acts as a regularizer encouraging modes having the same meta-mode to be close to each other.
@@
$\mathcal{L}_{\text{MWTA}}=\frac{1}{|D|}\sum^{|D|}_{i=1}\frac{1}{\pi_{\theta_{k^{*}}}}\sum^{K^{\prime}}_{k^{\prime}=1}-\text{log}\left(\pi_{\theta_{k^{*},k^{\prime}}}\times\mbox{Laplace}(\boldsymbol{\mu}_{\theta_{k^{*},k^{\prime}}},\mathbf{b}_{\theta_{k^{*},k^{\prime}}})\right)$ (10)
In addition, we add to both terms a classification loss such as the Cross-Entropy loss to optimize the mixture component weights. We provide more details on the HWTA loss in Appendix 0.C.
@@ 4.3 Hierarchical Light Transformer Ensemble
Trajectory forecasting models increasingly rely on Transformer architectures [41] due to their ability to capture spatial and temporal correlations effectively. Ensemble methods have shown promise in enhancing the performance of these DNNs. However, like in other domains, the growing size of models poses computational challenges when ensembling. To address this, we propose leveraging the over-parameterization inherent in DNNs [12] to construct ensembles of lighter Transformer-based architectures while preserving their representational capacity comparable to larger models. A crucial consideration in ensembling trajectory forecasting DNNs is the linear increase in parameters with the number of ensemble members. Packed Ensemble techniques [22] have been demonstrated to be an efficient ensembling of light Convolutional Neural Networks. Our focus is on adapting these principles for Transformer architectures, addressing the challenge of efficiently ensembling these models. In particular, we pioneered the introduction of the Grouped Multi-head Attention Layer. Using our loss HWTA, we introduce two ensemble-based variants of our techniques: HT-Ens, a straightforward ensemble of DNNs for trajectory prediction trained with HWTA, and HLT-Ens: a lightweight version of HT-Ens that we developed using the Grouped Multi-head Attention Layer.
@@
To encapsulate several smaller fully-connected layers into a larger one, we use Grouped Fully-Connected (GFC) layers. This concept draws inspiration from the idea that grouped convolutions [20] enable the simultaneous training of multiple branches in parallel, as demonstrated in prior works [46, 22]. The idea is to map a portion of the input to a portion of the output without any overlap.
@@
Figure 2: Block diagonal weight matrix for a Grouped Fully-Connected layer with $G=2$ groups (Note: only the non-zero weights are stored in memory).
Figure 3: Grouped Multi-head Attention layer operations for $H=2$ and $G=3$. Step 1 computes the query, key, and value features for each head using grouped fully-connected layers to ensure independence between all three groups. Step 2 depicts attention mechanisms for each group in each head. Step 3 illustrates the projection of the concatenation of the heads using a grouped fully-connected layer.
@@
Consider $Q\in\mathbb{R}^{n\times d_{q}}$ a query matrix, corresponding to $n\in\mathbb{N}$ query vectors of dimension $d_{q}\in\mathbb{N}$. Using $n_{v}\in\mathbb{N}$ key-value pairs with dimensions $d_{v}\in\mathbb{N}$, denoted as $K\in\mathbb{R}^{n_{v}\times d_{q}}$ and $V\in\mathbb{R}^{n_{v}\times d_{v}}$ respectively, the attention function $\text{Att}(Q,K,V)$ is defined as follows: $\text{Att}(Q,K,V)=(QK)^{T}V$.
@@
We want to create a grouped version of the multi-head attention function, thus ensuring independence between subnetworks. Similarly to the GFC layer, we assume the embedding dimensions $d$ of $Q$, $K$, and $V$ are divisible by $G\times H$. For the following, we will only detail the operations $Q$ as they are exactly the same for $K$ and $V$. With $Q=\text{concat}(Q_{1},...,Q_{G})$, we define the $G\times H$ learnable parameters for the heads, i.e., $W^{Q}_{g,h}\in\mathbb{R}^{d_{G}\times d^{H}_{G}}$ where $g\in\llbracket 1,G\rrbracket$, $d_{G}=d/G$, and $d^{H}_{G}=d/(GH)$. We apply our previously defined GFC operation on $Q$ for each head $h$ to obtain $Q^{\prime}_{h}$:
@@
Note that here, $W^{O}$ is the resulting block diagonal matrix from a GFC layer tensor weights as defined in Eq. 15. In our architecture, the number of groups corresponds to the number of subnetworks $M$ in our ensemble. As stacking grouped operations until the end ensures the independence between gradients, we will have independent subnetworks as long as other operations, such as Layer Normalization [2], can be grouped (e.g., Group Normalization [45] instead of Layer Normalization) or have independent actions on each element in the embedding dimension.
@@
By replacing fully connected and multi-head attention layers with their grouped counterpart, we reduce the size of the overall architecture. It comes down to the parameter reduction done in the GFC layer. In a classic fully connected layer, the weight matrix encapsulates $D_{j}\times D_{j+1}$ parameters. On the other hand, GFC has $G$ weight matrices containing $D_{j}\times D_{j+1}\times G^{-2}$ parameters, meaning that divide the number of parameters by the number of groups $G$. Hence, if we were to use groups to fit subnetworks within a DNN backbone, the size of the ensemble would decrease as its number of members increases. To tackle this problem, we introduce a parameter $\alpha\in\mathbb{R}$ to compensate for the effect of using groups. It can be interpreted as a width expansion factor affecting the embedding dimension size. This enables a more flexible control of the ensemble size. The sensitivity analysis corresponding to this hyper-parameter can be found in Appendix 0.E.
@@ 5 Experiments
This section reports our experiments to validate the properties of our proposed contributions to trajectory forecasting tasks. Our evaluation focuses on assessing the accuracy of the predictions and the uncertainty estimates associated with the distribution of the resulting mixture of predictions. We provide a performance comparison of two Transformer-based backbones to showcase the benefits of both our loss and efficient ensembling architecture on trajectory forecasting tasks. Training procedures are the same for all methods sharing the same backbone, we report the hyperparameters in Appendix 0.B.
5.1 Trajectory Forecasting
Table 1: Performance comparison (averaged over five runs) on Interaction and Argoverse 1 using AutoBots and ADAPT backbones. All ensembles have $M=3$ subnetworks and are followed by a KMeans algorithm to form $6$ trajectory clusters from which we take the centroids; we highlight the best performances in bold. For our method, we consider $\alpha=1.5$ for AutoBots network and $\alpha=2$ for ADAPT network. The number of parameters is expressed in millions. M.A. corresponds to the inference cost, i.e., the number of Mega multiply-add operations for a forward pass, which is estimated with the python package torchinfo [39].
Method $\text{mADE}_{1}$downarrow $\text{mADE}_{3}$downarrow $\text{mADE}_{6}$downarrow $\text{mFDE}_{1}$downarrow $\text{mFDE}_{3}$downarrow $\text{mFDE}_{6}$downarrow b-$\text{mFDE}_{6}$downarrow $\text{NLL}_{3}$downarrow $\text{NLL}_{6}$downarrow Acc $\uparrow$ #Prm $\downarrow$ M.A. $\downarrow$
Interaction AutoBots Single
Legacy 0.67 0.41 0.27 1.95 1.11 0.66 1.30 -33.70 -60.18 23.12
WTA 0.65 0.40 0.28 1.92 1.12 0.70 1.38 -33.18 -68.64 19.94
$\varepsilon$-WTA 0.65 0.40 0.28 1.88 1.12 0.70 1.36 -44.72 -68.23 21.54
EWTA 0.59 0.39 0.28 1.72 1.11 0.70 1.38 -34.85 -65.99 20.24
HWTA (Ours) 0.57 0.35 0.26 1.69 0.97 0.68 1.36 -52.65 -69.50 19.42 1.2 15.22
Ensemble
DE 0.63 0.38 0.27 1.86 1.03 0.67 1.29 -41.23 -50.93 30.06 3.6 45.66
HT-Ens (Ours) 0.56 0.33 0.25 1.66 0.92 0.63 1.25 -54.65 -61.97 27.05 3.6 45.66
HLT-Ens (Ours) 0.60 0.35 0.27 1.78 0.97 0.68 1.31 -51.69 -58.03 26.92 0.9 11.62
ADAPT Single
Legacy 0.71 0.43 0.30 1.97 1.15 0.73 1.37 -30.43 -55.14 23.32
WTA 0.71 0.45 0.31 1.97 1.18 0.75 1.41 -27.00 -53.94 22.46
$\varepsilon$-WTA 0.74 0.45 0.31 2.05 1.20 0.75 1.42 -28.61 -52.93 21.03
EWTA 0.64 0.43 0.33 1.82 1.16 0.82 1.50 -33.54 -50.99 23.33
HWTA (Ours) 0.63 0.40 0.30 1.78 1.08 0.73 1.35 -26.92 -53.11 23.18 1.4 1.27
Ensemble
DE 0.70 0.39 0.29 1.94 1.02 0.70 1.32 -41.37 -50.27 26.37 4.2 3.81
HT-Ens (Ours) 0.63 0.36 0.28 1.79 0.95 0.69 1.32 -47.2 -55.3 27.17 4.2 3.81
HLT-Ens (Ours) 0.56 0.37 0.29 1.62 1.00 0.72 1.31 -45.34 -54.69 32.25 2.0 0.86
Argoverse 1 AutoBots Single
Legacy 1.65 1.07 0.77 3.66 2.07 1.18 1.83 57.15 33.03 19.11
WTA 1.85 1.08 0.78 4.09 2.14 1.25 1.93 52.41 27.00 17.16
$\varepsilon$-WTA 1.76 1.08 0.80 3.87 2.16 1.30 1.98 50.45 27.58 16.75
EWTA 2.22 1.11 0.81 4.87 2.16 1.26 1.95 51.03 32.26 16.66 17.24
HWTA (Ours) ($K^{\star}=3$,$K=9$) 1.55 0.99 0.78 3.44 1.91 1.22 1.89 43.33 30.41 20.10 1.5 20.27
Ensemble
DE 1.59 0.99 0.77 3.54 1.88 1.21 1.85 44.35 35.57 27.24 4.4 51.71
DE($K=9$) 1.62 0.98 0.81 3.59 1.85 1.35 1.95 44.09 37.82 29.79 4.5 60.80
HT-Ens (Ours) ($K^{\star}=3$,$K=9$) 1.54 0.95 0.76 3.43 1.82 1.19 1.85 39.28 30.26 24.60 4.5 60.80
HLT-Ens (Ours) ($K^{\star}=3$,$K=9$) 1.57 0.96 0.78 3.50 1.82 1.27 1.88 43.00 35.67 29.75 1.1 15.48
ADAPT Single
Legacy 1.89 1.15 0.85 4.27 2.30 1.40 2.08 58.02 34.39 18.29
WTA 1.84 1.13 0.83 4.13 2.29 1.38 2.07 56.65 33.14 18.57
$\varepsilon$-WTA 1.89 1.11 0.83 4.23 2.22 1.38 2.07 55.56 33.34 18.94
EWTA 1.80 1.15 0.84 4.00 2.33 1.40 2.09 58.74 34.40 18.49
HWTA (Ours) 1.61 1.07 0.84 3.58 2.13 1.37 2.05 55.10 35.20 19.91 1.4 1.27
Ensemble
DE 1.98 1.09 0.82 4.42 2.15 1.36 2.02 46.67 37.32 23.87 4.2 3.81
HT-Ens (Ours) 1.83 1.02 0.81 4.07 1.99 1.31 1.97 42.67 34.98 24.21 4.2 3.81
HLT-Ens (Ours) 1.54 1.00 0.78 3.40 1.94 1.24 1.88 40.97 31.12 23.69 2.0 0.86
@@ 5.1 Trajectory Forecasting
We evaluate our method’s performance in a single agent setting on Argoverse v1.1 [8] and Interaction [48]. We provide additional experiments in Appendix 0.H. Argoverse 1 Motion Forecasting is a dataset that allows us to train a model on a vast collection of $323,557$ scenarios of $5$ seconds each. Each scenario is a 2D bird-eye-view of the scene, tracking the sampled object at $10$Hz. The Agoverse v1.1 challenge is to predict $3$ seconds trajectory forecasting using $2$ seconds of the past trajectory. The Interaction dataset contains road traffic data splitter from $11$ locations recorded using drones or fixed cameras. The recordings are split into scenarios of $4$ seconds at a $10$Hz frequency. The Interaction challenge proposes to predict $3$ seconds of motion trajectory using $1$ second of prior observations.
@@
To conduct our experiments on the Argoverse 1 and Interaction datasets, we leverage two light Transformer-based backbones: AutoBots [13] and ADAPT [1]. First, we compare our loss, HWTA, to the classical WTA, the relaxed WTA ($\varepsilon$-WTA), and the Evolving WTA (EWTA) losses as well as the original losses used to train the backbones we are using (we denote these as the Legacy loss). All methods set the number of modes to $6$ ($K=6$), and our approach considers $2$ meta-modes ($K^{\star}=2$ and $K^{\prime}=3$) unless otherwise indicated. Then, to demonstrate the possibility of reducing the number of parameters in an ensemble of Transformer models, we report the performance of baseline ensembles (DE), ensembling of hierarchical transformers (HT-Ens), and our hierarchical light ensembling (HLT-Ens). We post-process the outputs of methods producing more than $6$ modes using the K-Means algorithm on the final predicted positions to get $6$ trajectories.
@@
[…]
@@
Tab. 1 presents the average performance on the Argoverse 1 and the Interaction datasets over five runs. Using baseline architectures, HWTA consistently achieves attractive results compared to other optimization strategies on most metrics, especially in terms of $\text{mADE}_{1}$, $\text{mADE}_{3}$, $\text{mFDE}_{1}$, and $\text{mFDE}_{3}$. It demonstrates the ability of our loss to focus on high-probability density areas while retaining enough diversity to perform comparably with other methods. Concerning ensembling approaches, it is interesting to see that their benefits in terms of forecasting accuracy are most visible on $\text{mADE}_{k}$ and $\text{mFDE}_{k}$ with $k\in\{3,6\}$. Yet, they consistently improve the accuracy (Acc) of the models demonstrating the enhancement of mixture weights as the most confident mode is more often matched with the best one.
@@
Concerning HLT-Ens, its application to ADAPT outperforms the baseline ensemble counterpart by a large margin on $\text{mADE}_{1}$ and $\text{mFDE}_{1}$, and achieves similar results on other metrics, using a small fraction of the parameters and number of operations required by the other ensembles. The application to AutoBots provides also good results compared to DE, with much fewer parameters, but in contrast, seems to underperform slightly compared to HT-Ens. We argue that it might be the symptom of a representation capacity too small to be subdivided into $3$ networks. In Appendix 0.E, we discuss the influence of the representation capacity, and in Appendix 0.G, we provide more insight into the differences between HLT-Ens and DE in the context of trajectory forecasting.
@@
Tab. 2 demonstrates an improved stability induced by our loss in the optimization process. As WTA losses focus on the best prediction, the performance on the $\text{mADE}_{6}$ and $\text{mFDE}_{6}$ are consistent, i.e., low standard deviation across runs. However, we observe a higher performance variation on both $\text{mADE}_{1}$ and $\text{mFDE}_{1}$ suggesting instabilities in the optimization process. Concerning HWTA, in contrast, the standard deviation is much lower demonstrating the ability of our new loss to achieve consistent results across training runs.

### v3

@@ 4.2 Hierarchical Multimode Loss
Many existing MDN architectures rely on techniques similar to clustering, which can introduce challenges regarding robustness to changes in mode number. In contrast, our proposed approach embraces a hierarchical structure, enabling more flexibility. To integrate this hierarchical structure into our prediction mixture, we introduce a novel loss function aimed at softly enforcing proximity among modes belonging to the same meta-mode, termed the Hierarchical Winner-Takes-All (HWTA) loss. The HWTA loss, denoted as $\mathcal{L}$, comprises two distinct components: the meta-mode loss ($\mathcal{L}_{\text{meta}}$) and the Meta-Winner-Takes-All (MWTA) loss ($\mathcal{L_{\text{MWTA}}}$).
@@
with a hyperparameter $\gamma\in\left[0,1\right]$ to control the importance of both terms. We conduct a sensitivity study on this parameter in Appendix D. The meta-mode loss is the NLL of the meta-mixture, considering $K^{\star}\in\mathbb{N}$ as the number of meta-modes:
$\mathcal{L}_{\text{meta}}=$ $\frac{1}{|D|}\sum^{|D|}_{i=1}\sum^{K^{\star}}_{k^{\star}=1}-\text{log}\left(\pi_{\theta_{k^{\star}}}(X_{i})\right)$
$-\log\left(\mbox{Laplace}(Y_{i}|\ \overline{\boldsymbol{\mu}}_{\theta_{k^{\star}}}(X_{i}),\overline{\mathbf{b}}_{\theta_{k^{\star}}}(X_{i}))\right)$ (9)
This term enables backpropagating gradients to all modes, preventing hypothesis collapse and maximizing the likelihood of the meta-mixture.
@@
$\mathcal{L}_{\text{MWTA}}=$ $\frac{1}{|D|}\sum^{|D|}_{i=1}\frac{1}{\pi_{\theta_{k^{*}}}}\sum^{K^{\prime}}_{k^{\prime}=1}-\text{log}\big{(}\pi_{\theta_{k^{*},k^{\prime}}}\big{)}$
$-\log\big{(}\mbox{Laplace}(Y_{i}|\ \boldsymbol{\mu}_{\theta_{k^{*},k^{\prime}}},\mathbf{b}_{\theta_{k^{*},k^{\prime}}})\big{)}$ (10)
This term encourages modes having the same meta-mode to be close to each other as they will be optimized on the same subset of the data. In addition, we add to both terms a classification loss such as the Cross-Entropy loss to optimize the mixture component weights. We provide more details on the HWTA loss in Appendix C.
@@ 4.3 Hierarchical Light Transformer Ensemble
Trajectory forecasting models increasingly rely on Transformer architectures [42] for their ability to effectively capture spatial and temporal correlations. Ensemble methods have shown promise in enhancing the performance of these DNNs. However, like in other domains, the growing size of models poses computational challenges when ensembling. To address this, we propose leveraging the over-parameterization inherent in DNNs [13] to construct ensembles of lighter Transformer-based architectures while preserving their representational capacity comparable to larger models. A crucial consideration in ensembling trajectory forecasting DNNs is the linear parameter increase with the number of ensemble members. Packed Ensemble techniques [23] have been demonstrated to be an efficient ensembling of light Convolutional Neural Networks. We focus on adapting these principles for Transformer architectures, addressing the challenge of efficiently ensembling these models. In particular, we propose the introduction of the Grouped Multi-head Attention Layer. Using our loss HWTA, we introduce two ensemble-based variants of our techniques: HT-Ens, a Deep Ensemble of DNNs for trajectory prediction trained with HWTA, and HLT-Ens: a lightweight version of HT-Ens that we developed using the Grouped Multi-head Attention Layer.
@@
To encapsulate several smaller fully connected layers into a larger one, we use Grouped Fully-Connected (GFC) layers. This concept draws inspiration from the idea that grouped convolutions [21] enable the simultaneous training of multiple branches in parallel, as demonstrated in prior works [47, 23]. The idea is to map a portion of the input to a part of the output without overlap.
@@
Figure 2: Block diagonal weight matrix for a Grouped Fully-Connected layer with $G=2$ groups.
Figure 3: Grouped Multi-head Attention layer operations for $H=2$ and $G=3$. Step 1 computes the query, key, and value features for each head using grouped fully-connected layers to ensure independence between all three groups. Step 2 depicts attention mechanisms for each group in each head. Step 3 illustrates the projection of the concatenation of the heads using a grouped fully connected layer.
@@
Consider $Q\in\mathbb{R}^{n\times d_{q}}$ a query matrix, corresponding to $n\in\mathbb{N}$ query vectors of dimension $d_{q}\in\mathbb{N}$. Using $n_{v}\in\mathbb{N}$ key-value pairs with dimensions $d_{v}\in\mathbb{N}$, denoted as $K\in\mathbb{R}^{n_{v}\times d_{q}}$ and $V\in\mathbb{R}^{n_{v}\times d_{v}}$ respectively, the attention function is defined as follows: $\text{Att}(Q,K,V)=\text{softmax}\left(\frac{QK^{T}}{\sqrt{d_{q}}}\right)V$.
@@
We want to create a grouped version of the multi-head attention function, thus ensuring independence between subnetworks. Similarly to the GFC layer, we assume the embedding dimensions $d$ of $Q$, $K$, and $V$ are divisible by $G\times H$. For the following, we will only detail the operations $Q$ as they are precisely the same for $K$ and $V$. With $Q=\text{concat}(Q_{1},...,Q_{G})$, we define the $G\times H$ learnable parameters for the heads, i.e., $W^{Q}_{g,h}\in\mathbb{R}^{d_{G}\times d^{H}_{G}}$ where $g\in\llbracket 1,G\rrbracket$, $d_{G}=d/G$, and $d^{H}_{G}=d/(GH)$. We apply our previously defined GFC operation on $Q$ for each head $h$ to obtain $Q^{\prime}_{h}$:
@@
Here, $W^{O}$ is the resulting block diagonal matrix from a GFC layer tensor weights as defined in Eq. 15. In our architecture, the number of groups corresponds to the number of subnetworks $M$ in our ensemble. As stacking grouped operations until the end ensures the independence between gradients, we will have independent subnetworks as long as other operations, such as Layer Normalization [2], can be grouped (e.g., Group Normalization [46] instead of Layer Normalization) or have independent actions on each element in the embedding dimension.
@@
By replacing fully connected and multi-head attention layers with their grouped counterpart, we reduce the size of the overall architecture. It comes down to the parameter reduction done in the GFC layer. The weight matrix encapsulates $D_{j}\times D_{j+1}$ parameters in a classic fully connected layer. On the other hand, GFC has $G$ weight matrices containing $D_{j}\times D_{j+1}\times G^{-2}$ parameters, meaning that we divide the number of parameters by the number of groups $G$. Hence, if we were to use groups to fit subnetworks within a DNN backbone, the size of the ensemble would decrease as its number of members increases. To tackle this problem, we introduce a parameter $\alpha\in\mathbb{R}$ to compensate for the effect of using groups. It can be interpreted as a width expansion factor affecting the embedding dimension size. This enables a more flexible control of the ensemble size. The sensitivity analysis corresponding to this hyper-parameter can be found in Appendix E.
@@ 5 Experiments
Method $\text{mADE}_{1}$downarrow $\text{mADE}_{6}$downarrow $\text{mFDE}_{1}$downarrow $\text{mFDE}_{6}$downarrow $\text{NLL}_{3}$downarrow $\text{NLL}_{6}$downarrow
Interaction AutoBots Legacy [14] 0.67 0.27 1.95 0.66 -33.70 -60.18
WTA[12, 37] 0.65 0.28 1.92 0.70 -33.18 -68.64
$\varepsilon$-WTA[37] 0.65 0.28 1.88 0.70 -44.72 -68.23
EWTA[29] 0.59 0.28 1.72 0.70 -34.85 -65.99
HWTA (Ours) 0.57 0.26 1.69 0.68 -52.65 -69.50
ADAPT Legacy[1] 0.71 0.30 1.97 0.73 -30.43 -55.14
WTA[12, 37] 0.71 0.31 1.97 0.75 -27.00 -53.94
$\varepsilon$-WTA[37] 0.74 0.31 2.05 0.75 -28.61 -52.93
EWTA[29] 0.64 0.33 1.82 0.82 -33.54 -50.99
HWTA (Ours) 0.63 0.30 1.78 0.73 -26.92 -53.11
Argoverse 1 AutoBots Legacy[14] 1.65 0.77 3.66 1.18 57.15 33.03
WTA[12, 37] 1.85 0.78 4.09 1.25 52.41 27.00
$\varepsilon$-WTA[37] 1.76 0.80 3.87 1.30 50.45 27.58
EWTA[29] 2.22 0.81 4.87 1.26 51.03 32.26
HWTA (Ours) 1.52 0.78 3.38 1.25 70.05 32.27
ADAPT Legacy[1] 1.89 0.85 4.27 1.40 58.02 34.39
WTA[12, 37] 1.84 0.83 4.13 1.38 56.65 33.14
$\varepsilon$-WTA[37] 1.89 0.83 4.23 1.38 55.56 33.34
EWTA[29] 1.80 0.84 4.00 1.40 58.74 34.40
HWTA (Ours) 1.61 0.84 3.58 1.37 55.10 35.20
Table 1: Loss performance comparison (averaged over $5$ runs) on Interaction and Argoverse 1 using AutoBots and ADAPT models. We highlight the best performances in bold.
Strategy $\text{mADE}_{1}$downarrow $\text{mADE}_{6}$downarrow $\text{mFDE}_{1}$downarrow $\text{mFDE}_{6}$downarrow #Prm $\downarrow$ MAC $\downarrow$ Time (h) $\downarrow$
Interaction AutoBots top-$6$ 0.68 0.37 1.97 0.99 3.6 45.66 1.3
RIP[30] 0.68 0.40 1.96 1.09 3.6 45.66 1.3
Multipath++[41] 0.64 0.32 1.89 0.85 3.6 45.66 1.3
TENET, QCNet [43, 50] 0.64 0.27 1.89 0.67 3.6 45.66 1.3
HT-Ens. (Ours) 0.57 0.25 1.69 0.62 3.6 45.66 1.3
HLT-Ens. (Ours) 0.61 0.27 1.80 0.69 0.9 11.62 0.6
ADAPT top-$6$ 0.71 0.36 1.98 0.92 4.2 3.81 2.0
RIP[30] 0.75 0.37 1.84 0.96 4.2 3.81 2.0
Multipath++[41] 0.66 0.37 1.84 0.96 4.2 3.81 2.0
TENET, QCNet [43, 50] 0.70 0.29 1.94 0.70 4.2 3.81 2.0
HT-Ens. (Ours) 0.63 0.28 1.79 0.69 4.2 3.81 2.0
HLT-Ens. (Ours) 0.56 0.29 1.62 0.72 2.0 0.86 0.7
Argoverse 1 AutoBots top-$6$ 1.63 0.99 3.61 1.84 4.5 60.80 3.3
RIP[30] 1.60 1.05 3.56 1.98 4.5 60.80 3.3
Multipath++[41] 1.56 0.84 3.46 1.42 4.5 60.80 3.3
TENET, QCNet [43, 50] 1.62 0.81 3.59 1.35 4.5 60.80 3.3
HT-Ens. (Ours) 1.54 0.76 3.43 1.19 4.5 60.80 3.3
HLT-Ens. (Ours) 1.57 0.78 3.50 1.27 1.1 15.48 2.0
ADAPT top-$6$ 1.83 0.95 4.04 1.72 4.2 3.81 5.1
RIP[30] 1.64 1.07 3.68 2.03 4.2 3.81 5.1
Multipath++ [41] 1.62 0.98 3.61 1.84 4.2 3.81 5.1
TENET, QCNet [43, 50] 1.76 0.82 3.92 1.36 4.2 3.81 5.1
HT-Ens. (Ours) 1.72 0.81 3.84 1.31 4.2 3.81 5.4
HLT-Ens. (Ours) 1.54 0.78 3.40 1.24 2.0 0.86 2.3
Table 2: Ensemble strategies comparison (averaged over five runs) on Interaction and Argoverse 1 using AutoBots and ADAPT backbones. We highlight the best performances in bold.
@@
This section reports our experiments to validate the properties of our proposed contributions to trajectory forecasting tasks. Our evaluation focuses on assessing the accuracy of the predictions and the uncertainty estimates associated with the distribution of the resulting mixture of forecasts. We provide a performance comparison of four Transformer-based backbones to showcase the benefits of both our loss and efficient ensembling architecture on trajectory forecasting tasks. Training procedures are the same for all methods sharing the same backbone, and we report the hyperparameters in Appendix B. The results of Wayformer and SceneTransformer are detailed in Appendix H.
@@
We evaluate our method’s performance in a single agent setting on Argoverse 1 [8] and Interaction [49]. Argoverse 1 Motion Forecasting is a dataset that allows us to train a model on a vast collection of $323,557$ scenarios of $5$ seconds each. Each scenario is a 2D bird-eye-view of the scene, tracking the sampled object at $10$Hz. The Agoverse 1 challenge is to predict $3$ seconds trajectory forecasting using $2$ seconds of the past trajectory. The Interaction dataset contains road traffic data from $11$ locations recorded using drones or fixed cameras. The recordings are split into scenarios of $4$ seconds at a $10$Hz frequency. The Interaction challenge proposes to predict $3$ seconds of motion trajectory using $1$ second of prior observations.
@@
To conduct our experiments on the Argoverse 1 and Interaction datasets, we leverage four Transformer-based backbones: AutoBots [14], ADAPT [1] in Tab. 1, and SceneTransformer [35] and Wayformer [34] in Appendix H. First, we compare our loss, HWTA, to the classical WTA [37], the relaxed WTA ($\varepsilon$-WTA)[37], and the Evolving WTA (EWTA)[29] losses as well as the original losses used to train the backbones we are using (we denote these as the Legacy loss). All methods set the number of modes to $6$ ($K=6$), and our approach considers $2$ meta-modes ($K^{\star}=2$ and $K^{\prime}=3$) unless otherwise indicated in Appendix B. Then, to demonstrate the possibility of reducing the number of parameters in an ensemble of Transformer models, we report the performance of baseline ensembles (DE) with different forecast aggregation strategies: taking the top-$k$ most confident predictions, applying RIP [30], the EM algorithm as in [41], and K-Means [43, 50, 34]. We compare them to the ensembling of hierarchical transformers (HT-Ens) and our hierarchical light ensembling (HLT-Ens). We post-process the outputs of methods producing more than $6$ modes using the K-Means algorithm on the final predicted positions to get $6$ trajectories.
@@
We evaluate our models using standard metrics in the field [8, 6, 45, 49]: minimum Average Displacement Error ($\text{mADE}_{k}$) and minimum Final Displacement Error ($\text{mFDE}_{k}$). Subscript $k$ denotes that these metrics are calculated over the top-$k$ most confident forecasted trajectories. We define the best trajectory as having the closest endpoint to the ground truth over the $k$ trajectories. $\text{mADE}_{k}$ measures the average position-wise $\ell_{2}$ distance between the ground-truth and the best predicted trajectory. $\text{mFDE}_{k}$ corresponds to the $\ell_{2}$ distance between the ground-truth endpoint and the best trajectory one. In addition, we propose to study the $\text{NLL}_{k}$ with $k\in\{3,6\}$. We report the number of parameters (#Prm) expressed in millions, the number of Mega multiply-add operations for a forward pass (MAC), and the training Time to measure ensembling approach complexity.
(a) (left) Sub-modes of three meta-modes (distinct color). (right) meta-modes computed as the average of their sub-modes.
Figure 4: Examples of forecasts of an AutoBots model trained using the HWTA loss on the Argoverse 1 dataset. It has three meta-modes with two sub-modes each. The observed trajectory of the agent of interest is depicted with a dark blue solid line, while the grey dashed lines are the centerlines of road lanes. We represent the model’s forecasts in cyan, magenta, and yellow. Solid lines are meta-modes and dashed lines are sub-modes.
Quantitative Results:
@@
Tab. 1 presents the average models’ performance on the Argoverse 1 and the Interaction datasets over five runs. Using baseline architectures, HWTA consistently achieves attractive results compared to other optimization strategies on most metrics, especially in $\text{mADE}_{1}$ and $\text{mFDE}_{1}$. It demonstrates the ability of our loss to focus on high-probability density areas while retaining enough diversity to perform comparably with other methods.
Tab. 2 presents the average performance of models on the Argoverse 1 and Interaction datasets over five runs using different ensemble strategies. It is shown that HT-Ens, representing a Deep Ensembles with our HWTA loss, achieves state-of-the-art performance. Furthermore, HLT-Ens, which employs our strategy to reduce the computational budget of HT-Ens, matches HT-Ens’s performance while cutting the computational time by at least half. This demonstrates the significant advantages of HLT-Ens over other strategies.
@@
On a more general note, the K-Means algorithm as an aggregation strategy produces satisfactory results for diversity. Thus, we used it to aggregate the predictions of our ensembling methods: HT-Ens and HLT-Ens. Concerning HLT-Ens, its application to ADAPT outperforms the baseline ensemble counterpart by a large margin on $\text{mADE}_{1}$ and $\text{mFDE}_{1}$. It achieves similar results on other metrics, using a small fraction of the parameters and number of operations required by classical ensembles. The application to AutoBots also provides good results compared to DE, with much fewer parameters, but in contrast, it seems to underperform slightly compared to HT-Ens. We argue that it might be the symptom of a representation capacity too small to be subdivided into $3$ networks. In Appendix E, we discuss the influence of the representation capacity. In Appendix G, we provide more insight into the differences between HLT-Ens and DE in trajectory forecasting.
@@
Tab. 3 demonstrates an improved stability induced by our loss in the optimization process. As WTA losses focus on the best prediction, the performance on the $\text{mADE}_{6}$ and $\text{mFDE}_{6}$ are consistent, i.e., low standard deviation across runs. However, we observe a higher performance variation on both $\text{mADE}_{1}$ and $\text{mFDE}_{1}$, suggesting instabilities in the optimization process. Concerning HWTA, in contrast, the standard deviation is much lower, demonstrating our new loss’s ability to achieve consistent results across training runs.